# Data Contamination Issues in Brain-to-Text Decoding

### v2

@@ Abstract
Decoding non-invasive cognitive signals to natural language has long been the goal of building practical brain-computer interfaces (BCIs). Recent major milestones have successfully decoded cognitive signals like functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG) into text under open vocabulary setting. However, how to split the datasets for training, validating, and testing in brain-to-text decoding still remains controversial. In this paper, we conduct systematic analysis on current dataset splitting methods and find data contamination largely exaggerates model performance. Specifically, first we find the leakage of test subjects’ cognitive signals corrupts the training of a robust encoder. Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize seen information in test set. To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive dataset (e.g. fMRI, EEG). We also test the performance of SOTA brain-to-text decoding models under the proposed dataset splitting paradigm as baselines for further research.
11footnotetext: Work was done when the author was an intern in JD.com.22footnotetext: Corresponding author.
@@ 1 Introduction
Brain-computer interface (BCI) builds connections between human brain and external devices (e.g. computer). It has been widely researched in the field of neuroscience and has gained remarkable success like repairing damaged sight or restoring movement of disabled people (Polikov et al., 2005; Hochberg et al., 2012; Bouton et al., 2016). However, it’s still challenging in decoding cognitive signals to corresponding natural language when subjects read or hear text stimuli, especially for non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG) that are noisy and of low resolution (Mridha et al., 2021).
Figure 1: General frameworks of brain-to-text decoding and possible situations of data contamination.
Recent methods (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023) typically viewed brain-to-text decoding as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015) and adopted an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language. Specifically, as shown in Figure1, the encoder usually consists of a spatial and time series feature extractor. It can be trained either in an end-to-end manner with decoder (Figure 1 a) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 b). Despite recent success in model design, it still remains controversial in how to split the dataset for training, validating, and testing (Xi et al., 2023). Addressing this issue is urgent and meaningful, as fair evaluation of models is impossible without a widely recognized dataset splitting paradigm.
A cognitive dataset is usually formatted in signal-sentence pair. In most cases for brain-to-text decoding task, each sentence belongs to a certain task, so signal-sentence pair can be further divided into signal-task and task-sentence pair. Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task. However, all these methods suffer from data contamination on encoder side, decoder side, or both. As shown in Figure 1, to encoder if subjects’ cognitive signals in test set are mixed into training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals. As to decoder, situation gets worse if text stimuli are leaked. Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder is able to predict next token regardless of encoded cognitive signal representations.
@@
To address the above-mentioned problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides. We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the given format. In our method, the dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation set and test set won’t appear in training set, which means the trained encoder can’t get access to any brain information belonging to subjects in test set. (2) Text stimuli in validation set and test set won’t appear in training set. The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen text.
@@
  • We investigate current dataset splitting methods and analyze their influence on popular frameworks in brain-to-text decoding.
  • We prove the existence of data contamination in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
  • We propose the first splitting method without data contamination on public cognitive datasets. We also release a fair benchmark to evaluate different models’ generalization performance for further research in this domain.
Figure 2: Different cognitive dataset splitting methods.
@@ 2 Related Work
Cognitive Signal Cognitive signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue. Due to the high cost and complexity of invasive and partially invasive methods, it’s hard to apply them in building generic and practical BCIs. In this paper, we mainly focus on non-invasive signals EEG and fMRI. EEG signal is electrogram of the spontaneous electrical activity of the brain. Its frequencies usually range from 1 to 30 Hz, divided into several groups like alpha (4-13 Hz), beta (13-30 Hz), delta (0.5-4 Hz), theta (4-7 Hz). EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly. fMRI measures brain activity by detecting changes of blood flow. Blood flow of a specific region increases when this brain area is in use. The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004). Unlike EEG which samples brain signals continuously, fMRI samples according to the smallest time period named TR, usually at second level.
Brain-to-text Decoding Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021). These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between cognitive signals and words in vocabulary. Despite relatively good accuracy, these methods fail to generalize to unseen words. Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework, or use less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019). However, these models struggle to generate accurate and fluent sentences limited by decoder ability. Wang and Ji (2022) introduced the first open vocabulary EEG-to-text decoding model by leveraging the power of pre-trained language models. Xi et al. (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
@@ 3 Methodology
In this section, we will first introduce the definition of decoding cognitive signals to text and the general description of dataset format. Then we systematically analyze current dataset splitting methods and point out that all existing methods suffer from two kinds of data contamination: cognitive signal leakage and text stimuli leakage. Finally, a new dataset splitting method is proposed to avoid the above-mentioned two kinds of data contamination.
@@ 3.1 Task Definition
Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, $L$ consistent sentence segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Here $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$. To EEG, since signals corresponding to a complete sentence are available, we bond sentence (aka text stimuli) $T_{j}$ and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most circumstances, each sentence $T_{j}$ belongs to one certain task $M_{k}$. So the signal-sentence pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
@@
Figure 3: The process of our proposed dataset splitting method.
@@
In brain-to-text decoding, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021). As a result, if cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signal $F_{i}$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. So the dataset splitting rules can be formally defined as:
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{val},S_{test}\}$ (1)
$S_{val}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{train},S_{test}\}$ (2)
$S_{test}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{train},S_{val}\}$ (3)
3.2 Dataset Splitting Methods
@@
Current dataset splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, the five dataset splitting methods are characterised as (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2. Figure 2 vividly displays the differences between current dataset splitting methods. For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences. The line connecting two symbols indicates they are related to one sample. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s one sample where subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our dataset splitting method. As the splitting of validation set is similar to test set, we only consider training set and test set in this section.
@@
for training set. Method (b) splits the dataset according to tasks, which can be described as
@@
for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames
@@
However, there are slight differences between these methods. Method (c) views all the cognitive signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then union all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames while method (e) picks consecutive signal frames.
@@
Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR).
@@
The goal of brain-to-text decoding models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received text stimuli are new to the trained model. In this sense, we define two kinds of data contamination: cognitive signal leakage and text stimuli leakage. The data contamination situation of different methods is reflected in Figure 2. If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set. Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation. Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two cognitive signals too. Since fMRI signals need to be sampled continuously with a certain length $L$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L-1$ continuous part following. In this sense, for EEG cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage. The situation of method (b) is opposite to that of method (a), where there’s no text stimuli leakage but cognitive signal leakage. Method (c) and method (d) are similar. They suffer from both cognitive signal leakage and text stimuli leakage. Method (e) is in the same situation as method (b). For fMRI, method (c), (d), and (e) which seem the same for EEG are actually different splitting ways. The situation of data contamination for different methods is similar to EEG, except for method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples.
3.3 Our Method

### v3

# Cross-Subject Data Splitting for Brain-to-Text Decoding

@@ Abstract
Recent major milestones have successfully decoded non-invasive brain signals (e.g. functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG)) into natural language. Despite the progress in model design, how to split the datasets for training, validating, and testing still remains a matter of debate. Most of the prior researches applied subject-specific data splitting, where the decoding model is trained and evaluated per subject. Such splitting method poses challenges to the utilization efficiency of dataset as well as the generalization of models. In this study, we propose a cross-subject data splitting criterion for brain-to-text decoding on various types of cognitive dataset (fMRI, EEG), aiming to maximize dataset utilization and improve model generalization. We undertake a comprehensive analysis on existing cross-subject data splitting strategies and prove that all these methods suffer from data leakage, namely the leakage of test data to training set, which significantly leads to overfitting and overestimation of decoding models. The proposed cross-subject splitting method successfully addresses the data leakage problem and we re-evaluate some SOTA brain-to-text decoding models as baselines for further research.
@@ 1 Introduction
Brain-to-text decoding aims to recover natural language from brain signals stimulated by corresponding speech. Recent studies (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023; Duan et al., 2024) have successfully decoded non-invasive brain signals (e.g. fMRI, EEG) to text by applying deep computational neural networks. However, no consensus has reached on how to split the cognitive dataset for training, validating, and testing. Most of the prior work (Ye et al., 2023; Tang et al., 2023) performed subject-specific data splitting for training and evaluating decoding models. Under this splitting rule, data for training, validating, and testing all comes from one specific subject from the cognitive dataset. For example, Tang et al. (2023) picked three subjects out of seven and conducted model training and evaluation on the three subjects respectively. This subject-specific splitting method causes two main problems. First, it only utilizes a tiny part of the whole dataset. Since the collection of brain signals is costly and time-consuming, such splitting method results in significant waste of data resources. Second, it leads to the poor generalization of decoding models. As every subject’s brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects. Moreover, decoding models trained from scratch on limited data are prone to facing the overfitting problem.
Figure 1: Illustration of naturalistic language comprehension dataset for brain-to-text decoding. Path with the same color indicates one sample for training/validating/testing.
Some studies (Wang and Ji, 2022; Xi et al., 2023) began to shed light on cross-subject data splitting, which views all the subjects’ data as a whole and performs splitting according to a given ratio (e.g. 8:1:1 for the training set, validation set, and test set). Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024). However, unlike datasets for brain-to-image decoding, where subjects are guided to see different and unrepeated pictures, different subjects will hear the same story in the naturalistic language comprehension dataset for brain-to-text decoding, which challenges cross-subject data splitting. As shown in Figure 1, such dataset is usually formatted in subject-task-text-signal ($S$-$M$-$T$-$F$) pair, indicating the brain signal $F$ of subject $S$ stimulated by hearing text $T$ from task $M$. Current cross-subject data splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized as five categories: (1) split by subjects $S$, (2) split by tasks $M$, (3) split by randomly picking signal frames $F$, (4) split by randomly picking signal frames under certain task $F$-$M$, (5) split by randomly picking consecutive signal frames under certain task $F$-$M$. However, based on our observations, all these splitting methods suffer from data leakage problem, namely part of the test data is mixed into the training set, which leads to overfitting in model training and overestimation in model evaluation.
Specifically, modern brain-to-text decoding models follow an encoder-decoder manner. We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) for investigating the damage of data leakage. The former is an end-to-end encoder-decoder framework, while the latter first pre-trains the encoder and then applies it in the decoder training. Experiments support that data leakage affects model training on both the encoder side and decoder side. For the encoder component, if subjects’ brain signals in the test set are mixed into the training set, the encoder will become overfitting and fail to well represent unseen subjects’ brain signals. As to the decoder, the situation gets worse if text stimuli are leaked. Since the decoder follows an auto-regressive manner and generates token by token, data leakage will cause the decoder to memorize seen paragraphs during the teacher-forcing training stage, which leads to poor generalization to unseen text.
@@
To avoid data leakage and fairly evaluate the performance of brain-to-text decoding models, we propose a cross-subject data splitting criterion. We focus on fMRI and EEG signals in this study, although the proposed criterion could be applied to any datasets satisfying the prescribed format. In the proposed method, the dataset is split according to subject-text ($S$-$T$) pairs with the following rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set. (2) Text stimuli in validation set and test set will not appear in training set. The decoder learns to reconstruct language with brain signals instead of memorizing seen text.
@@
  • To the best of our knowledge, we propose the first cross-subject data splitting criterion for brain-to-text decoding.
  • We comprehensively analyze current cross-subject data splitting methods and find that all existing methods suffer from data leakage problem, which severely affects the training and evaluation of decoding models.
  • Some SOTA brain-to-text decoding models are re-evaluated under the proposed cross-subject data splitting method as baselines for further research.
Figure 2: Different splitting methods for cognitive dataset. (Color printing is preferred.)
@@ 2 Related Work
Brain Signal
Brain signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue. In this paper, we mainly focus on non-invasive signals EEG and fMRI. EEG signal is electrogram of the spontaneous electrical activity of the brain, with frequencies ranging from 1 Hz to 30 Hz. EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly. fMRI measures brain activity by detecting changes of blood flow. Blood flow of a specific region increases when this brain area is in use. The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004). Unlike EEG which samples brain signals continuously, fMRI samples based on a fixed time interval named TR, usually at second level.
Brain-to-text Decoding
Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021). These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between brain signals and words in vocabulary. Despite relatively good accuracy, these methods fail to generalize to unseen words. Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework or using less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019). However, these models struggle to generate accurate and fluent sentences limited by decoder ability. Wang and Ji (2022) introduced the first open vocabulary EEG-to-text decoding model by leveraging the power of pre-trained language models. Xi et al. (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
@@ 3 Methodology
In this section, we will first introduce the formal definition of brain-to-text decoding and the general description of dataset format. Then we systematically analyze current cross-subject data splitting methods and point out that all existing methods suffer from two kinds of data leakage issues: brain signal leakage and text stimuli leakage. Finally, a cross-subject splitting criterion is proposed which avoids the above-mentioned data leakage problems.
@@ 3.1 Task Definition
Given the brain signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, consistent text segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$, where $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$, and $|T_{j}|=|F_{ij}|=L$. To EEG, since signals corresponding to complete text stimuli are available and they are continuous, we bond text $T_{j}$ (i.e. text stimuli) and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most scenarios, each text stimulus $T_{j}$ belongs to one certain task $M_{k}$. So the signal-text pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
@@
Figure 3: The process of our proposed cross-subject data splitting method. (Color printing is preferred.)
@@
One of the purposes of cross-subject splitting is to endow models the ability to decode unseen subject’s brain signal. As a result, if brain signal $F_{ij}$ appears in test set $S_{test}$, any signal $F_{i*}$ belonging to subject $i$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. The above splitting rules for training set can be formulated by Cartesian product:
$S_{train}=F_{train}\times T_{train},$ (1)
where
$F_{train}=\{F_{ij}|i\in I\},$ (2)
$I=\{i|F_{ij}\notin S_{val},S_{test},\forall j\},$ (3)
and
$T_{train}=\{T_{kj}|T_{kj}\notin S_{val},S_{test}\}.$ (4)
Such rules are also applicable to validation and test set. We omit their displays here for simplicity.
3.2 Current Splitting Methods
@@
Current cross-subject data splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, the five dataset splitting methods are characterised as (1) split by subjects $S_{i}$, (2) split by tasks $M_{k}$, (3) split by randomly picking signal frames $F_{ij}$, (4) split by randomly picking signal frames under certain task $F_{ij}$-$M_{k}$, (5) split by randomly picking consecutive signal frames under certain task $F_{ij}$-$M_{k}$, corresponding to (a), (b), (c), (d), (e) in Figure 2. Figure 2 vividly displays the differences between current dataset splitting methods. In this example, we choose 4 subjects ($S_{1}$ to $S_{4}$) with 3 tasks ($M_{1}$ to $M_{3}$) each containing 4 ($T_{11}$ to $T_{14}$), 3 ($T_{21}$ to $T_{23}$), 4 ($T_{31}$ to $T_{34}$) text stimuli respectively. $F_{ij}$ indicates the brain signal of $i$-th subject stimulated by $j$-th text under task $M_{k}$, e.g. $F_{21}$ means brain signal of $S_{2}$ hearing $T_{11}$.
The line connecting two symbols indicates they are related to one sample in dataset. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it indicates that subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our data splitting method. As the splitting of validation set is the same as test set, we only consider training set and test set in this section for simplicity.
@@
for training set. Method (b) splits the dataset according to tasks, which is described as
@@
for training set. Method (c), (d), and (e) all split the dataset according to brain signal frames
@@
However, there are slight differences between these three methods. Method (c) views all the brain signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then unions all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames, while method (e) picks consecutive signal frames.
@@
Table 1: Results of Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR).
@@
We first point out the data leakage problems in current splitting methods through the analysis of training set and test set composition. Specifically, following the definition in subsection 3.1, two kinds of data leakage, brain signal leakage and text stimuli leakage, are defined. The data leakage situation of different methods can be reflected through visualization in Figure 2. Lines between $S_{i}$ and $M_{k}$ indicate brain signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation. If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set. Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two kinds of brain signal too. Since fMRI signals need to be sampled continuously with a certain length $L$, one sample shown in Figure 2 (e.g. $S_{1}$-$M_{1}$-$T_{11}$-$F_{11}$) is actually the first part of one fMRI sample, with $L-1$ continuous brain signal frames following (e.g. $F_{12},F_{13},\ldots$). In this sense, brain signal leakage doesn’t exist in method (a) for EEG, but method (a) suffers from text stimuli leakage. Text stimuli leakage does not exist in method (b) but brain signal leakage exists. Method (c) suffers from both brain signal leakage and text stimuli leakage. Method (d) and (e) are the same to EEG, with brain signal leakage. For fMRI, the situation of data leakage for different methods is similar to EEG, except for method (d) and (e), which are the same for EEG but actually different for fMRI. Method (d) suffers from both brain signal leakage and text stimuli leakage while in method (e) text stimuli leakage happens in the overlap between training samples and test samples.
3.3 Cross-Subject Splitting Criterion